OpenStudy (anonymous):

Line m contains points (1, -3) and (2, 2). Which of the following pairs of points define a line parallel to line m?

OpenStudy (rulnick):

Slope = (2-(-3)) / (2-1) = 5. So any line of slope 5 will be parallel.

OpenStudy (anonymous):

(1, 1) and (6, 2) (0, 0) and (-1, 1) (-4, 0) and (5, 5) (0, 0) and (1, 5)

OpenStudy (luigi0210):

Find the slope of the points using \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (rulnick):

Still don't see? It would be the last one, slope (5-0)/(1-0) = 5.

OpenStudy (rulnick):

In Luigi0210's formula for slope, (x1,y1) is the first pair and (x2,y2) is the second. See it now?